DVILJK(1) DVILJK(1) NAME dvilj4,dvilj4l,dvilj2p,dvilj−convertaTeXDVIfiletoPCL,forHPLaserJetprinters SYNOPSIS dvilj[−−Dnumber][−ccopies][−dmode][−D][−E][−eoutputfile][−ffrompage][−g][−hheaderfile][−l][−Mflag][−mmag][−ppages][−q][−r][−ttopage][−spagesize][−v][−Vmode][−w] [−xxoff(mm)][−yyoff(mm)][−Xxoff(dots)][−Yyoff(dots)][−z][−][filename] DESCRIPTION DviljandsiblingsconvertTeX-output.dvifilesintoPCL(theHewlett-PackardPrinterControl Language)commandssuitableforprintingonaHPLaserJet+,HPLaserJetIIP(usingdvilj2p), HPLaserJet4(usingdvilj4),andfullycompatibleprinters. Dvilj4alsoprovidessupportforthebuiltinIntellifontandTrueTypefonts;TFMandVFfilesare providedinthedistribution.VirtualfontsarenotsupporteddirectlybyDvilj,soyoumustrun dvicopy(1)first,orotherwiseexpandthevirtualfonts Allprogramsaregeneratedfromonesourcefiledvi2xx.c.Theyareusuallycalledviaashell scriptorasalineprinterfilter.Onesuchscript,dvihp,isincludedinthedistribution.Itprovides command-lineoptionssimilartothoseofdvips(1). OPTIONS --DnumberDebugaccordingtothebitsofnumber;seetheKpathseamanualforvalues.Also turnsongeneralverbosity. -cnumberprinteachpagenumbertimes(includingoriginal) -dnumberprintinduplexmode.Theoption-d1meanslongedgebinding,-d2standsforshort edgebinding.Thisoptionisonlyselectablefromdvilj2p,andisknowntoworkonthe LaserJetIId,LaserJetIIId,andLaserJet4. -D[-]Double-sidedprinting-seebelow;-D1printsoddsidesonly,-D2isforevenpages,a trailing-suppressesprintingofemptypaddingpages. -ESwitchprinterintoeconomode(dvilj4only). -eoutfileSendoutputtooutfileinsteadofthebasenameofdvifileextendedwith‘.lj’.Ifoutfile is‘-’(asin‘-e-’)theoutputisdirectedtostandardoutputandcanbedirectlypiped intoaspooler. -fpagenumPrintpageswithTeXpagenumbersgreaterthanorequaltopagenum. -g‘‘go’’mode:donotresetprinteratstartofjob. -hheaderfile CopyheaderfiletooutputbeforetranslationoftheDVIfile.Usedtosendarbitrary commandstothelaserprinter. -lprintoutputinlandscapemode;supportedonlybydvilj2panddvilj4. -MflagDo(not)callmktexpktogeneratefontsondemand.-M1and-Mareequivalent. -M0enablesinvokingmktexpk.(Thisisthesameasdvips.)Thedefaultbehavior (i.e.,ifno-Moptionisgiven)ischosenatcompile-time;asdistributed,mktexpkis enabled.-Misonlyvalidifdvi2xxwascompiledwithsupportformktexpk. -mmagnification Specifythemagnificationtouse.Thefollowingareavailable: -m#usermage.g.,#900 -m0equivalentto-m#1000 -mhequivalentto-m#1095 -m1equivalentto-m#1200 -mqequivalentto-m#1250 -m2equivalentto-m#1440 -m3equivalentto-m#1728 10June1996 1 DVILJK(1) DVILJK(1) -m4equivalentto-m#2074 -m5equivalentto-m#2488 -ppagecount printpagecountpagesafterthestartingpage. -qQuietoperation:omitfinalstatistics,etc. -rtogglepagereversal;bydefault,pagesarereversed. -spagesizeSpecifythepagesize: -s1:executive(7.25inx10.5in) -s2:letter(8.5inx11in) -s3:legal(8.5inx14in) -s26:A4(210mmx297mm) -s80:monarch(3.875inx7.5in) -s81:commercial-10(4.125inx9.5in) -s90:internationalDL(110mmx220mm) -s91:internationalC5(162mmx229mm) AllvaluesarefromtheLJIItechnicalreferencemanual.Ifthe-scommandline optionisomitted,noclippingisperformedatthebottomofthepage. -tpagenumPrintpageswithTeXpagenumberslessthanorequaltopagenum. -vVerbosemode:listfontsused,etc. -Vmodecompatibilitymode.UseKforKyocera,BforBrother,6forLJ6printers. -wOmitwarnings. -xoffsetglobalx-offsetinmmonoutputpagerelativeto1inch/1inchorigin -yoffsetglobaly-offsetinmmonoutputpagerelativeto1inch/1inchorigin -Xoffsetglobalx-offsetindotsonoutputpage(overrides1inch/1inchorigin) -Yoffsetglobaly-offsetindotsonoutputpage(overrides1inch/1inchorigin) -zprintatestpagewiththepagecounteratendofjob -Takeinputfromstandardinput(mustbeseekable),writetostandardoutput. Ifnoargumentsarespecifiedabriefusagemessageisdisplayed. \specialCOMMANDS TeX‘\special’commandsthatareimplemented: \special{file=<filename>}copyspecifiedfiletotheprinter;eithervector-orbitmap-graphicsPMPcommands \special{hpfile=<filename>}Sameas‘file’. others: ThereareafewotherspecialcommandsimplentedfortheLJ-series,whichshouldonlybeusedin connectionwiththedistributedstylefiles(e.g.,graybox.sty).Theymaychangeinforthcoming versions. YoucanincludeEncapsulatedPostScript(EPS)filesinthePCLfilesoutputbydviljifyouconvertthemtoPCLfirst.Thedistributionincludesascript‘epspcl’todothis(usingGhostscript,so yourGhostscriptmusthavethenecessaryLaserJetdriver(s)compiledin). DOUBLEPAGEPRINTING Thecommandlineoption-Dcausestheprintertoprintindoublepagemode,thatis,roughly speaking,printingfirsttheevenpagesandthentheoddpagesofadocument.If-Disusedto printbothevenandoddpages(-Disusedwithouttheoptional1or2)thefollowingwillhappen afterhalfofthepagesisprinted:LJ:theprinterstops,showsmessage"FEED"andswitcheson manualfeedled.Theprintercontinueswheneithertheprintedpagesarefedintothemanualpage 10June1996 2 DVILJK(1) DVILJK(1) feeder,ortheoperatorpushesthecontinuebutton(afterplacingfirsttheprintedpagesintothe papertray). If-D1wasspecifiedonlyoddpageswillbeprinted,-D2canbeusedforevenpages.Atrailing‘-’ (asin-D-,-D1-or-D2-)supressesprintingofemptypageswhicharegeneratedforpaddingother-wise. Thedriverallowsthe-Doptionforbothprint-directions(front-to-backorback-to-frontprinting). Ittriestobecleverincertaincases:whenadocumenthasanoddnumberofpagesitinsertsin therightplaceanemptypagewhenadocumentstartswithpage0thedriverinsertsanempty pageafterthe0-pagewhenapieceofadocumentisprintedthatstartswithanevenpage-number itprintsthispageatthebacksideofthefirstpagetoensurethattheoddsidesarealwaysonthe frontside.The0-pageistreatedlikeanoddpage. ThedoublesideoptionworksforallLJs,notonlyforthenewDmodelwithdoesdoubleside-print-ing(actually,Idon’tknowwhetheritmakesmuchsenseonthenewLJIIDatall). Caveatonthedoubleside-option:Whenitisusedfrequentlyoronverylargefiles,itmightbecome necessarytocleantheprintermuchmorefrequently. Thecommandlineoption-pXcanbeusedtoabortprintingaftertheoutputofXpages(canlead tostrangeresultstogetherwith-D).Ipersonallydislikethisoption.Thesecurewayofprintinga partofadocumentisusing-f(from)and-t(to). ENVIRONMENT DviljkusesthesameenvironmentvariablesandalgorithmsforfindingfontfilesasTeXandits friendsdo.SeethedocumentationfortheKpathsealibraryfordetails.(Repeatingithereistoo cumbersome.) DVILJFONTS,ifset,overridesallotherfontpaths. DVILJMAKEPK,ifset,overrides‘mktexpk’asthenameofthecommandtoexecutetocreatea PKfileifoneisn’tfound. TheDVILJSIZESvariablemaybesettoindicatewhichsizesoffontsareavailable.Itshouldconsistofalistofnumbersseparatedbycolons.Ifthelistbeginswithacolon,thesystemdefault sizesareused,aswell.Sizesareexpressedindotsperinch;decimalsmaybeusedfor‘‘pxl’’files: forexample,a300dotsperinchfilemagnifiedbyhalfastepcomesoutto1643dotsperfive inches,whichshouldbeencodedas328.6.dviljtriestheactualsizeofthefontbeforetryingany ofthegivensizes. KPATHSEADEBUG:TraceKpathsealookups;setto-1forcompletetracing. AUTHOR GustafNeumann,WirtschaftsuniversitaetWien neumann@wu-wien.ac.at dviljkchangesbykb@mail.tug.org;emailbugreportstotex−k@mail.tug.org. 10June1996 3